### v1

# Water vapor emission at the warm cavity wall of the HD 100546 disk as revealed by ALMA

Abstract
We present spatially resolved ALMA observations of the water line at 183 GHz in the disk around the Herbig star HD 100546. The water vapor emission peaks at the inner edge of the warm dust cavity, located au from the central star. We attribute this to thermal desorption at the water snowline, shifted outward at the dust cavity wall directly heated by the intense radiation. This represents the first spatially resolved image of the water snowline using ALMA observations of the main water isotopologue in a protoplanetary disk. The water emission morphology peaking inside the first dust ring is consistent with previous ALMA detections of oxygen-bearing complex organic molecules in the disk, including thermally desorbed methanol. These findings signal that warm cavities of transition disks provide ideal targets to directly reconstruct the spatial distribution of water vapor and the snowline location with ALMA, and directly connect water vapor emission to ice desorption of complex organic species.
show]luna.rampinelli@unimi.it
1 Introduction
Water is of primary interest in astronomy, as it plays a fundamental role in the emergence of life on Earth, liquid water being an efficient solvent that facilitates the development of chemical complexity. Understanding how water was delivered to Earth is a central question in astronomy (A. Morbidelli et al., 2012). However, answering this question is challenging, as it is closely intertwined with the complex chemical and physical history of the building blocks of planets and their interplay with water in the natal environments, the protoplanetary disks (E. F. van Dishoeck et al., 2021).
Water observations in the protoplanetary disk stage are extremely valuable, as they offer a unique opportunity to observationally characterize the link between the atmospheric composition of the evolved exoplanetary population and their birthplace (A. J. Cridland et al., 2016). In addition, the water in these sources is likely formed before the central star did, thus injecting all systems with a similar water abundance at the time of their formation (J. J. Tobin et al., 2023; A. Andreu et al., 2023; M. Leemker et al., 2025b). Being a primary oxygen carrier, water plays a crucial role in determining the chemical make-up of the planet-forming material, by often dominating the C/O ratio in the gas and ice phases (K. I. Öberg et al. 2011; 2023; C. Eistrup et al. 2018). Beyond the water snowline, at a midplane temperature below K, the water freezes on dust grains, enriching the ice with oxygen sequestered from the gas phase. The location of the snowline not only regulates the chemical composition of the disk reservoir but also promotes planet formation by facilitating grain growth and coagulation (D. Schoonenberg & C. W. Ormel, 2017; J. Drążkowska et al., 2023).
Water observations in disks have been performed mainly with infrared campaigns, and have confirmed the main water reservoirs predicted by disk thermo-chemical models (see, e.g., K. M. Pontoppidan et al., 2010; A. Banzatti et al., 2017; E. F. van Dishoeck et al., 2021, and references therein). Water gas is expected in three main regions in protoplanetary disks: (1) inside the water snowline, typically within a few astronomical units from the central star. This is a warm reservoir of thermally desorbed water at temperatures above K, with high abundances typically (P. Woitke et al., 2009; E. F. van Dishoeck et al., 2014); (2) in the outer disk. This is the cold component ( K) photodesorbed from icy grains, at intermediate layers where UV photons can penetrate. The abundance can reach up to and it is regulated by photodissociation and water self-shielding (E. A. Bergin et al., 2010; M. R. Hogerheijde et al., 2011); (3) in the disk surface layer. This is the hot component reformed in the gas phase at K. The high temperature required to unlock the gas-phase reaction pushes this region to higher layers, where the abundance can reach up to as a result of the balance between photodissociation and self-shielding (T. Bethell & E. Bergin, 2009). We refer the reader to E. F. van Dishoeck et al. (2021); L. M. Pirovano et al. (2022) for a detailed description.
A first characterization of the water emission in protoplanetary disks was provided by IR campaigns with Spitzer, revealing correlations with stellar and disk properties, such as the stellar temperature and luminosity, accretion luminosity, disk dust mass and size, but also dynamical processes such as the inward drift of icy pebbles (see e.g. K. M. Pontoppidan et al., 2010; C. Salyk et al., 2011; J. R. Najita et al., 2013; A. Banzatti et al., 2020). Higher resolution water spectra of the inner disk of protoplanetary disks obtained with the JWST revealed multiple temperature contributions associated to different water reservoirs (A. Banzatti et al., 2023a; I. Kamp et al., 2023), with the component consistent with the water sublimation temperature becoming more prominent at longer wavelengths (, A. Banzatti et al., 2023b; M. Temmink et al., 2024). The cold water reservoir photodesorbed in the outer disk was detected by Hershel in the disks of TW Hya and HD 100546, revealing a weaker emission than previously predicted by models, as a result of the settling of icy grains over time (E. A. Bergin et al., 2010).
However, IR observations with current facilities lack the spatial resolution to resolve the water emission and directly constrain the snowline location, which is typically within only a few astronomical units for T Tauri stars and within au for Herbig stars (P. Woitke et al., 2009; S. Notsu et al., 2017). J. J. Tobin et al. (2023) showed the first direct constraint of the water snowline location (at au) from ALMA observations of heavier water isotopologues in the disk around the outbursting V883 Ori star. S. Facchini et al. (2024) showed the first spatially and spectrally resolved ALMA observations of the main water vapor isotopologue in the protoplanetary disk of HL Tau. These results demonstrate the potential of ALMA in revealing the content, distribution, and location of the snowline of water vapor in disks. However, even though less important than in the IR, dust absorption can occult mm-wave water line emission arising from the disk midplane. Further imaging of water transitions at mm wavelengths in protoplanetary disks with optically thin inner regions can circumvent this problem (M. Leemker et al., 2025a).
In this work we present the first spatially resolved ALMA observations of water vapor in a disk hosting a wide dust cavity. The source is described in Sect. 2, while observations used to perform the analysis, self-calibration and image reconstruction are outlined in Sect. 3. We discuss the main results in Sect 4, and summarize the main conclusions in Sect. 5.
2 The source
HD 100546 is a Herbig star (M M⊙, C. Walsh et al. 2017, L 23.5 L⊙, M. Vioque et al. 2018), at a distance of 110 pc (Gaia Collaboration 2016, 2023). Two dust rings at and au have been detected in ALMA continuum observations (C. Walsh et al., 2014; J. E. Pineda et al., 2019; D. Fedele et al., 2021), while scattered light observations revealed small- and large-scale spiral arm structures (A. Garufi et al., 2016; K. B. Follette et al., 2017; E. Sissa et al., 2018). Both wide orbit and close-in forming planet candidates have been suggested from sub-millimeter and infrared observations, asymmetric SO emission, and CO kinematic signatures (T. Currie et al., 2015; E. Sissa et al., 2018; S. Casassus & S. Pérez, 2019; J. E. Pineda et al., 2019; S. D. Brittain et al., 2019; S. Pérez et al., 2020; A. S. Booth et al., 2023).
HD 100546 shows a line-rich spectrum with a variety of molecules detected with ALMA, among which complex organic molecules (COMs) such as methanol and methyl formate, showing ring shape emission and/or azimuthal variations (A. S. Booth et al., 2023; L. Keyte et al., 2023; A. S. Booth et al., 2024; M. Leemker et al., 2024; L. Evans et al., 2025). In particular, methanol has an excitation temperature of K inside 110 au, which is consistent with thermal desorption origin (L. Evans et al., 2025). Thermally-desorbed water is therefore expected to be detected in the same region as methanol, as they have similar sublimation temperatures (V. Wakelam et al., 2017; M. Minissale et al., 2022). M. Leemker et al. (2024) modeled ALMA line emission observations of the HD 100546 disk with DALI (S. Bruderer et al., 2012; S. Bruderer, 2013), predicting the water snowline location to be at au, at the inner edge of the first dust ring.
Gas-phase water was previously detected in the disk around HD 100546 with Herschel (E. F. van Dishoeck et al., 2021; L. M. Pirovano et al., 2022). Both cold and warm lines with upper state energies ranging from to K were targeted, but only the two cold lines at 557 GHz (E K) and 1113 GHz (E K) were detected, and attributed to the cold photodesorbed water reservoir. The emitting region was inferred to be outside 40 au from the spectrally resolved line profiles (E. F. van Dishoeck et al., 2021).
L. M. Pirovano et al. (2022) modeled the Herschel observations of HD 100546 using DALI (S. Bruderer et al., 2012; S. Bruderer, 2013) and constrained the water abundance in the three reservoirs discussed in Sect. 1. Their fiducial model predicts low water abundances in the thermally desorbed component () and in the hot reservoir formed in the gas phase (), while the photodesorbed component at intermediate layers outside the dust ring at au needs to be relatively abundant () to reproduce the observations.
3 Observations
In this Letter, we analyzed ALMA band 5 observations of HD 100546 from project 2023.1.01431.S (PI S. Facchini), which covers the line at GHz. The observations consist of three Execution Blocks (EBs), obtained in C-5 configuration on the 2024-06-30 (EB0 and EB1) and 2024-07-02 (EB2), with a maximum baseline of 2.5 km. The time on source for each EB was 47 minutes, for a total of 2.35 hours in exquisite weather conditions corresponding to a Precipitable Water Vapor (PWV) ranging from 0.21 to 0.45 mm across the EBs. The spectral setup is made of two spectral windows (spw) centered on the water line at GHz and the line at GHz, respectively, and two continuum spws centered at 172.0 and 185.5 GHz with a 1.875 GHz bandwidth each. Each of the two line spws has a bandwidth of 117 MHz, and a native spectral resolution of 122 kHz ( km s) for the spw and 30.5 kHz (0.05 km s) for the spw. After the ALMA Pipeline cross-calibration, we self-calibrated the data following the Pipeline designed for the exoALMA Large Program (R. A. Loomis et al., 2025), and using the CASA software version 6.5.4 (CASA Team, 2022).
3.1 Self-calibration
After flagging lines within a km srange around the systemic velocity, we averaged the data into 125 MHz wide channels. We applied a first round of phase-only self-calibration to individual EBs, combining spws and scans on EB-long time intervals. We then performed a -plane alignment as described by R. A. Loomis et al. (2025). The flux rescaling was not performed at this stage of the self-calibration since phase decoherence was found in the data (see R. A. Loomis et al. 2025 for more details). We then concatenated all EBs and applied self-calibration iteratively on progressively shorter time intervals (EB-long, 360s, 120s, 60s, 18s, and 6s). Only spws and polarizations (no scans) were combined for time intervals shorter than 360s. Phase self-calibration was performed using models produced through tclean, cleaning down to 6. An elliptical mask was used during the cleaning, with Position Angle , a major axis of to include the faint outer ring at au, and a minor axis of , where is the disk inclination (A. S. Booth et al., 2024).
Phase self-calibration improved phase coherence and flux rescaling was not necessary since flux offsets among different EBs were within 4%. We therefore proceeded with two rounds of amplitude self-calibration, using EB-long and scan-long intervals, respectively, and combining spws and polarizations. For this step we cleaned down to . We then applied the gain solutions and phase shifts (in the original order) to the full dataset, including spectral lines. Finally, we binned the data into 30s intervals and performed the continuum subtraction using the uvcontsub task. We report a continuum flux of mJy inside the de-projected circular mask used during the self-calibration, a noise RMS of 21.5 Jy beam-1, and a peak S/N ratio of 2240 at the end of the self-calibration.
Particular care was taken to estimate the noise RMS in the water spw, because of the low transmission caused by the atmospheric telluric line of water. The spectrum of the RMS and its impact on the water spectrum is discussed in detail in Appendix A.
3.2 Imaging
The analyses presented in this work were applied to spectral cubes of the water line reconstructed through the CASA software (CASA Team, 2022) using the tclean task (J. A. Högbom, 1974; T. J. Cornwell, 2008). We used the multiscale deconvolver with scales [0,5,10,20,30] pixels, a pixel size of , and a flux threshold of . To place the CLEAN components we used the same elliptical mask built for the self-calibration and used to image the continuum.
To find the best balance between sensitivity and spatial resolution, we imaged the line with Briggs weighting spanning a range of robust parameters between (beam size = , RMS 8 mJy beam-1) and (beam size = , RMS 5 mJy beam-1). However, note that the RMS varies substantially in the spectral range of the water emission, see Appendix A for more details. Similarly, to find the best balance between spectral resolution and S/N, we imaged the cube with a channel width of km sand km s. For the results presented in this Letter we used cubes both with a channel width of and km s, and imaged with a robust parameter of 0.5 and 2.0. Figure 5 in Appendix B shows the channel maps of the water line spectral cube, imaged with natural weighting and a channel width of 1.2 km s. We applied a correction for the so-called JvM effect (S. Jorsater & G. A. van Moorsel, 1995) following I. Czekala et al. (2021), to correctly recover the flux of the water line.
3.3 Moment maps, radial profile and spectrum
We calculated the integrated intensity map showed in the middle panel of Fig. 1, from the cube imaged with natural weighting to maximize the sensitivity. We integrated the cube between km sand km s, after visual inspection of the data with no additional masking, using the bettermoments package (R. Teague & D. Foreman-Mackey, 2019). The water integrated intensity (moment 0) map is compared to the continuum one in Fig. 1. Most of the water emission originates from within the bright continuum ring at au, while there is a fainter azimuthally asymmetric component emitting from outside the continuum ring at au towards south/west. The right panel of Fig. 1 shows the intensity weighted velocity map, obtained by integrating in the same velocity range used to calculate the moment 0 map. The (moment 1) map shows the typical blue- and red-shifted pattern around the systemic velocity of 5.7 km s, expected for a disk in Keplerian rotation.
We calculated the water line flux to be mJy km sby spatially integrating the velocity integrated intensity map inside an elliptical mask with a major axis of 100 au and a minor axis of , using the disk geometrical parameters (PA = and ) from A. S. Booth et al. (2024). The uncertainty is evaluated as the standard deviation of the flux measured in the same elliptical mask used to evaluate the line flux, but outside the emitting area of the integrated intensity map. We took the maximum number of ellipses that was possible to place on the image without overlap, performing the procedure on images that are not primary beam corrected, to ensure uniform noise. The flux was extracted from the JvM-corrected cube to avoid overestimating the line flux, while the uncertainty was evaluated from the non-JvM corrected image to avoid underestimating the noise (L. Rampinelli et al., 2024).
From the integrated intensity map of the water emission we extracted the radial profile of the integrated intensity, shown in Fig. 2, using the GoFish package (R. Teague, 2019). The uncertainty on the radial profile was extracted as:
where is the standard deviation over the first five and last five channels of the cube used to extract the moment 0 map, is the channel width, is the number of channels over which we integrated to extract the moment 0, and is the number of beams in each annulus over which we calculated the radial profile of the integrated intensity (which is set to one if ).
We show the radial profile obtained from the cube imaged with a robust parameter of 0.5 in favor of a higher spatial resolution to highlight possible radial features. The inner peak of the water emission coincides with the inner edge of the dust cavity wall, as visible from the comparison of the solid blue and dashed black profiles in Fig. 2. The outer azimuthally asymmetric emission observed in the integrated intensity map in Fig. 1 is also visible in the radial profile in Fig. 2, which shows an outer emission bump peaking at au. We will further discuss the emission morphology of water vapor in Sect. 4.
We extracted the disk-integrated spectrum from the cube imaged with natural weighting, using the GoFish package (R. Teague, 2019). The water line spectrum obtained by integrating over elliptical mask used to extract the water flux is shown in the left panel of Fig. 3. The spectrum shows the typical double-horn feature expected for a Keplerian disk, and centered at the systemic velocity of km s, indicated by the vertical dashed red line. The S/N associated to the peak of the spectrum is 2.6. The uncertainty is 30 mJy and it is evaluated as standard deviation of the 88 signal-free velocity bins (blue errorbar in the top right corner of Fig. 3). We then extracted the spectrum from the same region but correcting for the Keplerian rotation of the disk (R. Teague, 2019), as shown in the right panel of Fig. 3. The spectral shifting boosts the peak S/N to 5.7, with the uncertainty decreasing to 16.5 mJy.
We highlight that the width of the spectrum in the right panel of Fig. 3 should not be interpreted as the water linewidth, since it is affected both by beam and spectral smearing due to the limited spatial (beam major axis 35 au) and spectral (channel width = 1.2 km s) resolution of the cube used to extract it. Moreover, additional sources of uncertainty possibly affecting the Keplerian shift are associated to the assumed geometrical parameters (inclination, PA, emitting surface), stellar mass, and distance of the source.
4 Discussion
The presented observations show the first spatially resolved detection of the main water vapor isotopologue with ALMA in a transition disk. As anticipated in Sect. 3.3, the water emission at 183.31 GHz shows an inner bright peak at au. In the following sections we discuss this result in light of the previous dust and line observations of the disk around HD 100546.
4.1 Water vapor at the warm cavity wall
The peak of the radial profile of the water integrated intensity in Fig. 2 is at the cavity wall of the dust disk, at a radius of () au. The astrometric accuracy on the peak of the profile in Fig. 2 is estimated as 111See Sect. 10.5.2 Astrometric Observations & Astrometric Accuracy in the ALMA Cycle 12 Technical Handbook, which implies a offset between the disk cavity wall at au and the peak of the dust ring at au. We suggest this emission originates from thermally desorbed water at the snowline.
The location of the water snowline is typically only a few astronomical units away from the central star for T Tauri stars and within au also for warmer Herbig disks, as the high water binding energy implies a high sublimation temperature of K. However, HD 100546 is a transition disk around a Herbig star, with an intrinsically warmer temperature structure (M. Leemker et al., 2022) and a central dust and gas cavity (J. E. Pineda et al., 2019; S. Pérez et al., 2020; L. Wölfer et al., 2023; A. S. Booth et al., 2024). These physical conditions result in an efficiently irradiated cavity wall reaching temperatures higher than the water sublimation temperature (M. Leemker et al., 2024), and pushing the water snowline to larger radii. This is also consistent with the thermo-chemical model from M. Leemker et al. (2024) predicting the water snowline to be located at au at the cavity wall.
While less severe than at IR wavelengths (M. Leemker et al., 2025a), optically thick dust can partially hide water vapor emission. The rapid drop outside au in the radial profile of the water integrated intensity in Fig. 2 could also be a consequence of the optically thick dust ring at au, instead of water freeze-out. However, high spatial resolution ALMA band 7 continuum observations show that the peak of the continuum brightness temperature (at au) is K, well below the water sublimation temperature of K, suggesting that the water snowline is located at the inner edge of the dust ring at au. This is also supported by radiative transfer models showing that the midplane dust temperature steeply drops below K at the cavity wall (J. E. Pineda et al., 2014, 2019; L. Keyte et al., 2023; M. Leemker et al., 2024).
The thermal desorption origin of the observed water emission at the dust cavity wall is supported by the fact that multiple oxygen-rich molecules including methanol have been suggested to thermally sublimate in the HD 100546 disk (A. S. Booth et al., 2024; L. Evans et al., 2025).
To summarize, the evidence outlined above points to the conclusion that the water snowline is located at the cavity edge of the dust disk at au in the HD 100546 disk.
4.2 Comparison with Herschel observations
Previous observations of water vapor in this source with Herschel only revealed cold water transitions ( and K), associated to the photodesorption layer outside au. Hot ( K) and warm ( K) water lines were also targeted with Herschel but not detected in the disk of HD 100546 (E. F. van Dishoeck et al., 2021; L. M. Pirovano et al., 2022).
The hot water lines with high K) are expected to trace the hot water reservoir produced in the gas phase in the upper layers and inside the dust cavity (see Sect. 1), where photodissociation limits the water abundance. Water self-shielding has been shown to mitigate this effect when the water column density is high ( cm-2, T. Bethell & E. Bergin 2009; A. D. Bosman et al. 2022; J. K. Calahan et al. 2022). However, HD 100546 is a transition disk with a large dust cavity which is also depleted in gas. The right panel of Fig 6 in Appendix C represents the radial profile of the water column density from the thermochemical model by M. Leemker et al. (2024), which shows that the water column density in the disk cavity is expected to be low enough to allow efficient photodissociation. We highlight that this model includes photodissociation only in the vertical direction, and not in the radial one. The non-detection of the hot Herschel lines with K which trace the hot water reservoir in the upper layers and inside the cavity can be thus explained by the efficient photodissociation (L. M. Pirovano et al., 2022).
On the other hand, the water column density is higher at the disk cavity edge at au where the water snowline is expected to be located (see the right panel of Fig. 6), thus favoring water self-shielding against photodissociation. The warm Herschel lines with K and the ALMA line with K, which are expected to trace closer to the water snowline, are therefore expected to be detected.
These warm Herschel lines could go undetected if they are sub-thermally excited, as their critical density is higher than the local gas density. The left panel of Fig. 6 shows the critical density of the Herschel lines presented by L. M. Pirovano et al. (2022) (solid lines) and of the ALMA 183 GHz line (dashed line), as a function of temperature, calculated following A. Faure et al. (2024), under the optically thin assumption, and neglecting absorption, stimulated emission, and background emission (see Appendix C). The critical densities of the warm Herschel lines with K are cm-3 at K, and approximately a factor of three higher than the ALMA line. However, we highlight that these warm Herschel lines have high Einstein coefficients , and are therefore expected to be optically thicker than the ALMA 183 GHz line with . In the optically thick regime, photon trapping can consistently lower the critical density (Y. L. Shirley, 2015). This, together with the fact that the gas density is expected to be cm-3 at the disk cavity wall (M. Leemker et al., 2024), suggests that sub-thermal excitation cannot explain the non-detection of the warm Herschel lines against the detection of the ALMA line.
The non-detection of warm water vapor with Herschel could be due to the high continuum optical depth of the inner dust ring at Herschel wavelengths with respect to ALMA, potentially hiding water emission in this region in the FIR. On the other hand, the optical depth of dust is much lower at ALMA wavelengths thus allowing to detect the warm 183 GHz ALMA line ( K) emitting at the dust cavity edge.
Finally, we report a marginal detection (peak in the channel maps) of azimuthally asymmetric water emission peaking at au in the West and South sides of the disk (see the 8.10 km schannel in Fig. 5). In this context, we note that HD 100546 has a dynamically active disk with large-scale spiral arms (A. Garufi et al., 2016). The faint bump is also visible in the radial profile of the integrated intensity (Fig. 2) but the S/N is too low () to robustly conclude on its significance. Interestingly, photodesorption of water ice has been proposed in the cold outer gap between the two dust rings at and au (M. Honda et al., 2016; L. M. Pirovano et al., 2022).
4.3 Water vapor emission and the molecular complexity of HD 100546
A. S. Booth et al. (2024) presented ALMA line emission observations of a rich sample of molecular tracers in the disk of HD 100546, most of them showing a double ringed emission similar to what is observed in ALMA continuum observations. M. Leemker et al. (2024) showed that chemistry rather than the underlying density structure is responsible for the observed molecular rings. In particular, oxygen and carbon rich molecules show the peak of the emission just inside or outside the inner dust ring at au, respectively. In addition, most of the molecules show a second peak just outside the outer dust ring at au (A. S. Booth et al., 2024). This morphology suggests a radially varying C/O ratio, higher at the dust rings, where and CN show bright emission (M. Leemker et al., 2024). This picture is supported by the detection of methanol and SO inside the inner dust cavity (L. Evans et al., 2025), and it is also consistent with the water morphology presented in this work. As shown in Fig. 2, the inner bright peak of water emission at au is just inside the inner dust ring, while the outer faint water emission peaking at au is within the gas gap, both anti-correlating with the dust rings where the C/O ratio is proposed to be above one (M. Leemker et al., 2024).
The observed water emission morphology is expected to anti-correlate with the emission, the latter being a major destruction pathway for water (T. G. Phillips et al., 1992; E. A. Bergin et al., 1998; M. Leemker et al., 2021). shows a centrally peaked emission and an outer ring co-located with the outer dust ring (A. S. Booth et al., 2024). The observed faint water emission at au is therefore consistent with a region of fainter emission. While the radial profile of the water integrated intensity in Fig. 2 shows marginal evidence of a decreasing trend inside au, the spatial resolution of both the and observations does not allow to robustly conclude about their anti-correlation inside the inner dust cavity. The marginal inner drop in could also be due to a drop in the gas density, as the HD 100546 shows an inner gas cavity. On the other hand, is a good tracer of the ionization rate and if it became optically thicker than water in the inner gas cavity, it would follow the drop in the gas density (M. Leemker et al., 2024). However, we highlight that the detection of water vapor is consistent with the low /CO column density ratio extracted for the HD 100546 disk (A. S. Booth et al., 2024), which could either result from a low ionization rate or the presence of gas-phase water.
5 Conclusions
In this work we present the first spatially resolved observations of the main water vapor isotopologue with ALMA in a transition disk. Our observations of the HD 100546 disk show:
1.
water emission peaking at the warm cavity wall of the transition disk. The high temperature in the dust cavity induced by the intense radiation field pushes the snowline to larger radii, allowing water to thermally desorb at the inner edge of the dust cavity;
2.
the first direct image of the water snowline in a protoplanetary disk from spatially resolved observations of the main water isotopologue;
3.
the gas phase distribution of water vapor, complementing previous observations of O-rich COMs in the transition disk. This reveals that planets forming inside the warm cavity of the transition disk are exposed to a water-rich chemistry which can leave a crucial imprint on the atmospheric composition of planets forming in the disk;
4.
the potential of ALMA of tracing the water vapor distribution in planet-forming disks thanks to the extremely favorable weather conditions reached at the observatory site, and the high resolution and sensitivity that can be achieved.
Appendix A Noise in the water spw
We report a spectral dependence of the noise RMS in the water spw. The left panel of Fig. 4 shows the noise RMS evaluated inside a circle with a radius of 500 au, as a function of velocity. The noise RMS was extracted from non-JvM corrected cubes (to avoid underestimating the noise) imaged with natural weighting and a channel with of 1.2 km s. The vertical dotted line shows the systemic velocity of 5.7 km s(in LSRK velocity frame), while the vertical dashed line corresponds to a velocity of 0 km sin the topocentric frame. The x-axes show the velocity in LSRK (bottom) and topocentric frame (top). The strong spectral dependence of the noise level is due to the lower transmission at the peak of the water telluric line, which is shifted with respect to the LSRK systemic velocity of the science target, peaking at km sin LSRK frame. As expected, the peak of the noise rms is correctly centered at km sin the topocentric frame, with the Lorentzian wings clearly recognizable throughout the spectral window. The observed velocity shift in the LSRK frame with respect to the topocentric frame at the location of ALMA is consistent with the observing dates and times, and with the source coordinates. In particular, the three EBs were taken on 2024-06-30 at 20:55 UTC, 2024-06-30 at 22:24 UTC, and 2024-07-02 at 21:07 UTC. The small difference in dates only induces a small shift in LSRK velocity.
EB0 and EB1 were taken in worse weather conditions than EB2, with the former showing a mean PWV of 0.4 mm, and the latter 0.2 mm. The noise RMS is significantly higher for EB0 and EB1 than for EB2 (see the left panel of Fig. 4), due to the exponential dependence with the atmospheric transmission within the telluric line. Due to the lower signal-to-noise ratio during the first two EBs, the bandpass calibration of the water spw on EB0 and EB1 was performed averaging over 10 and 6 velocity channels, respectively, while no averaging was applied to EB2. This has a direct consequence in the retrieved continuum-subtracted spectrum of the water spw. The right panel of Fig. 4 shows the spectrum of the water line, for the three EBs separately (EB0 and EB1 were shifted up for clarity), inside a circle with a 50 au radius. The spectra show a small dip (also visible in Fig. 3) at the peak of the spw rms. As expected from the different weather conditions and spectral averaging during bandpass calibration for EB0 and EB1, the spectrum from EB2 shows the smallest dip.
We highlight that this effect could be particularly important when scheduling water observations with ALMA (in particular for the 183 GHz water line), depending on the observing date and science target, since the trough in the atmospheric transmission could partially overlap with the water line of the astrophysical source of interest, thus resulting in both an elevated rms level and artificial absorption features.
Appendix B Water channel maps
Figure 5 shows the channel maps of the detected water line in the HD 100546 disk. The cube was imaged with natural weighting and a channel width of km s(see Sect. 3.2 for more details on the imaging procedure).
Appendix C Water observations with Herschel and ALMA
The left panel of Fig. 6 shows the critical density of the Herschel lines in L. M. Pirovano et al. (2022) (solid lines) and of the ALMA 183 GHz line (dashed line), as a function of temperature. The profiles are colored based on their upper state energy. The critical densities are calculated following A. Faure et al. (2024):
where is the Einstein coefficient of the spontaneous emission and is the collision rate out of the level into the level . The upward collision rate is related to the downward one as follow (Y. L. Shirley, 2015):
where is the degeneracy, is the energy, and is the kinetic temperature. The critical density is calculated under the assumption of optically thin water emission, and neglecting absorption, stimulated emission, and background emission.
The right panel of Fig. 6 shows the radial profile of the water column density obtained from the thermochemical model built by M. Leemker et al. (2024) using DALI (S. Bruderer et al., 2012; S. Bruderer, 2013) for the HD 100546 disk. We highlight that this model aimed at reproducing the emission morphology of a variety of molecules detected with ALMA, in particular CO isotopologues, HCN, CN, , NO, and , but it was not specifically built to reproduce the water emission and observations in this source.

### v2

@@
 show]luna.rampinelli@unimi.it
-1 Introduction
+I Introduction
 Water is of primary interest in astronomy, as it plays a fundamental role in the emergence of life on Earth, liquid water being an efficient solvent that facilitates the development of chemical complexity.
@@
 The source is described in Sect.
-2, while observations used to perform the analysis, self-calibration and image reconstruction are outlined in Sect.
-3.
-We discuss the main results in Sect 4, and summarize the main conclusions in Sect.
-5.
-2 The source
+II, while observations used to perform the analysis, self-calibration and image reconstruction are outlined in Sect.
+III.
+We discuss the main results in Sect IV, and summarize the main conclusions in Sect.
+V.
+II The source
 HD 100546 is a Herbig star (M M⊙, C.
@@
 Bruderer, 2013) and constrained the water abundance in the three reservoirs discussed in Sect.
-1.
+I.
 Their fiducial model predicts low water abundances in the thermally desorbed component () and in the hot reservoir formed in the gas phase (), while the photodesorbed component at intermediate layers outside the dust ring at au needs to be relatively abundant () to reproduce the observations.
-3 Observations
+III Observations
 In this Letter, we analyzed ALMA band 5 observations of HD 100546 from project 2023.1.01431.S (PI S.
@@
 Loomis et al., 2025), and using the CASA software version 6.5.4 (CASA Team, 2022).
-3.1 Self-calibration
+III.1 Self-calibration
 After flagging lines within a km srange around the systemic velocity, we averaged the data into 125 MHz wide channels.
@@
 The spectrum of the RMS and its impact on the water spectrum is discussed in detail in Appendix A.
-3.2 Imaging
+III.2 Imaging
 The analyses presented in this work were applied to spectral cubes of the water line reconstructed through the CASA software (CASA Team, 2022) using the tclean task (J.
@@
 (2021), to correctly recover the flux of the water line.
-3.3 Moment maps, radial profile and spectrum
+III.3 Moment maps, radial profile and spectrum
 We calculated the integrated intensity map showed in the middle panel of Fig.
@@
 We will further discuss the emission morphology of water vapor in Sect.
-4.
+IV.
 We extracted the disk-integrated spectrum from the cube imaged with natural weighting, using the GoFish package (R.
@@
 Moreover, additional sources of uncertainty possibly affecting the Keplerian shift are associated to the assumed geometrical parameters (inclination, PA, emitting surface), stellar mass, and distance of the source.
-4 Discussion
+IV Discussion
 The presented observations show the first spatially resolved detection of the main water vapor isotopologue with ALMA in a transition disk.
 As anticipated in Sect.
-3.3, the water emission at 183.31 GHz shows an inner bright peak at au.
+III.3, the water emission at 183.31 GHz shows an inner bright peak at au.
 In the following sections we discuss this result in light of the previous dust and line observations of the disk around HD 100546.
-4.1 Water vapor at the warm cavity wall
+IV.1 Water vapor at the warm cavity wall
 The peak of the radial profile of the water integrated intensity in Fig.
@@
 To summarize, the evidence outlined above points to the conclusion that the water snowline is located at the cavity edge of the dust disk at au in the HD 100546 disk.
-4.2 Comparison with Herschel observations
+IV.2 Comparison with Herschel observations
 Previous observations of water vapor in this source with Herschel only revealed cold water transitions ( and K), associated to the photodesorption layer outside au.
@@
 The hot water lines with high K) are expected to trace the hot water reservoir produced in the gas phase in the upper layers and inside the dust cavity (see Sect.
-1), where photodissociation limits the water abundance.
+I), where photodissociation limits the water abundance.
 Water self-shielding has been shown to mitigate this effect when the water column density is high ( cm-2, T.
@@
 Pirovano et al., 2022).
-4.3 Water vapor emission and the molecular complexity of HD 100546
+IV.3 Water vapor emission and the molecular complexity of HD 100546
 A.
@@
 Booth et al., 2024), which could either result from a low ionization rate or the presence of gas-phase water.
-5 Conclusions
+V Conclusions
 In this work we present the first spatially resolved observations of the main water vapor isotopologue with ALMA in a transition disk. Our observations of the HD 100546 disk show:
@@
 The cube was imaged with natural weighting and a channel width of km s(see Sect.
-3.2 for more details on the imaging procedure).
+III.2 for more details on the imaging procedure).
 Appendix C Water observations with Herschel and ALMA